# May I Ask a Follow-up Question? Understanding the Benefits of Conversations in Neural Network Explainability

### v1

@@ Abstract.
Research in explainable AI (XAI) aims to provide insights into the decision-making process of opaque AI models. To date, most XAI methods offer one-off and static explanations, which cannot cater to the diverse backgrounds and understanding levels of users. With this paper, we investigate if free-form conversations can enhance users’ comprehension of static explanations, improve acceptance and trust in the explanation methods, and facilitate human-AI collaboration. Participants are presented with static explanations, followed by a conversation with a human expert regarding the explanations. We measure the effect of the conversation on participants’ ability to choose, from three machine learning models, the most accurate one based on explanations and their self-reported comprehension, acceptance, and trust. Empirical results show that conversations significantly improve comprehension, acceptance, trust, and collaboration. Our findings highlight the importance of customized model explanations in the format of free-form conversations and provide insights for the future design of conversational explanations.
@@ 1. Introduction
The rapid advancement of Artificial Intelligence (AI) is powered by opaque deep neural networks (DNNs), which are difficult to interpret by humans (Bodria et al., 2023). The lack of transparency prevents verification of AI decisions by human domain experts and is especially concerning in areas of high-stake decisions, such as healthcare and law enforcement, where erroneous algorithmic decisions could lead to severe harm to human lives (Caruana et al., 2015; Cai et al., 2019) and erosion of public trust (Quinn et al., 2021; Powles and Hodson, 2017). To improve the interpretability (or explainability, which we use interchangeably in this paper) of AI models, numerous eXplainable Artificial Intelligence (XAI) methods have been proposed (see survey articles (Yang et al., 2019; Danilevsky et al., 2020; Bodria et al., 2023)). Perhaps surprisingly, despite the research effort, a number of recent studies (Poursabzi-Sangdeh et al., 2021; Zhang et al., 2020; Bansal et al., 2021; Wang and Yin, 2021; Lakkaraju et al., 2022; Liao et al., 2020) find that the explanations often do not resolve user confusion regarding the neural networks they are purported to explain. This curious phenomenon warrants further investigation.
@@
Having reviewed the literature, we attribute user dissatisfaction to three major reasons. First, the explanations do not properly account for layperson users’ knowledge of machine learning, which may not be sufficient to mentally establish the causal relation between the explanations and the model behaviors being explained (Poursabzi-Sangdeh et al., 2021; Ma et al., 2023; Springer and Whittaker, 2019). Communication theory posits that effective communication requires the experts to establish common ground with layperson users (Clark and Marshall, 1981; Clark and Brennan, 1991). However, experts usually find it hard to accurately estimate what laypeople know (Wittwer et al., 2008; Wilkesmann and Wilkesmann, 2011; Miller, 2019). To make matters worse, underestimating and overestimating the user’s knowledge level are equally detrimental to communication (Wittwer et al., 2008; Lakkaraju et al., 2022). As a result, it is little wonder that the explanations designed by experts are almost always at a mismatch with the user’s actual knowledge level.
Second, existing techniques are designed to explain only one facet of model behaviors, and follow-up questions naturally arise (Liao et al., 2020; Lakkaraju et al., 2022; Abdul et al., 2018), but existing interaction paradigms typically do not account for this. For example, feature attribution techniques (Sundararajan et al., 2017; Selvaraju et al., 2017; Simonyan et al., 2013; Ribeiro et al., 2016; Hu et al., 2018; Alvarez-Melis and Jaakkola, 2017) reveal the input features that the model rely on to make a prediction, but does not show how predictions would change if the input features change. This may be answered using counterfactual explanations (Verma et al., 2020; Wachter et al., 2017; Sharma et al., 2019; Karimi et al., 2020; Tran et al., 2021; Mothilal et al., 2020; Poyiadzi et al., 2020; Wijaya, 2021). However, current interaction paradigms of model explanation typically supply only a single static explanation, and assumes that explanation has been accomplished after that; they are not prepared for the possibility of follow-up questions.
@@
Finally, users of XAI have diverse backgrounds, intentions, and information needs (Suresh et al., 2021; Wang and Yin, 2021; Ehsan et al., 2021; Liao et al., 2020). For example, Liao and Varshney (Liao and Varshney, 2021) identifies five different objectives of users of explanations, including model debugging, assessing the capabilities of AI systems, making informed decisions, seeking recourse or contesting the AI, and auditing for legal or ethical compliance. Naturally, one static explanation is not sufficient to satisfy all objectives and purposes. Therefore, researchers (Lakkaraju et al., 2022; Abdul et al., 2018; Schmid and Wrede, 2022; Rohlfing et al., 2020; Dazeley et al., 2021; Cheng et al., 2019) have suggested injecting interactivity to model explanations in order to establish common ground, address knowledge gaps, and to create customized explanations that adapt to the users.
@@
Existing work on interactive explanations can be broadly categorized into two types. The first type, interactive machine learning (Amershi et al., 2014; Fails and Olsen Jr, 2003), generates model explanations that users rely on to provide feedback and suggestions to the machine learning model. Their primary goal is to improve machine learning performances of by interacting with machine learning practitioners, rather than explaining model behaviors to layperson users. In this interactive setting, explanations have been shown to improve user satisfaction (Smith-Renner et al., 2020) and feedback quality (Liang et al., 2020; Kulesza et al., 2015). The second type aims to demonstrate the input-output relationship by allowing users to freely modify input features and observe changes in the output, while showing current feature attribution explanations (Cheng et al., 2019; Liu et al., 2021a; Tenney et al., 2020; Hohman et al., 2019). This type of interactivity has been shown to improve user understanding (Cheng et al., 2019) and perceived usefulness (Liu et al., 2021a) of AI models. However, the effective use of these interactive approaches still requires a rudimentary understanding of achine learning, such as the generic relation between input and output, or what model properties the interpretations reveal. These interactive explanations still cannot answer most types of follow-up questions laypeople may have about the explanations.
@@
Free-form conversations that accompany static explanations are arguably the most versatile mode of interaction as they allow users to ask arbitrary follow-up questions and can give tailored explanations of neural networks for users with diverse backgrounds (Liao et al., 2020; Dazeley et al., 2021; Feldhus et al., 2022; Lakkaraju et al., 2022). Through interviews with decision-makers, Lakkaraju et al. (Lakkaraju et al., 2022) discover that they have a strong preference for explanations in natural language dialogue. They also outline five principles of interactive explanations and argue that conversational explanations fulfill all five. Hence, the authors argue that conversations are an appealing form of explanation for users with limited machine learning knowledge. With the progress in conversational characters (Zhang et al., 2022; Ni et al., 2023; Shuster et al., 2022), especially knowledge-based question answering (Lan et al., 2021; Luo et al., 2023; Zhang et al., 2023) powered by large language models (Touvron et al., 2023; Ouyang et al., 2022; Zhao et al., 2023), AI systems that can answer questions about their own decisions appear to be within our reach in the near future. However, before investing effort to develop such a chatbot, it is beneficial to empirically quantify the effects of conversational explanations.
@@
In the current study, we conduct Wizard-of-Oz experiments to investigate how conversations assist users in understanding static explanations, improving acceptance and trust in XAI methods, and selecting the best AI models based on explanations. Specifically, we first present participants with examples of static explanations and measure their objective understanding and subjective perception of static explanations. Afterward, the participants carry a conversation with an AI system through a textual interface and seek to clarify any doubts. An XAI expert acts as the mock AI system and attempts to answer the questions faithfully through the textual interface. The participants are motivated to understand the explanations as they need to select the best-performing machine learning model only using explanations when doing objective evaluation. After the conversation, participants complete the same pre-conversation measurements. By comparing results before and after the conversation, we can estimate the effects of conversations on static explanations.
@@
We design both an objective evaluation to measure users’ understanding of static explanations and a subjective evaluation to gauge users’ perceptions. The subjective evaluation can complement the objective evaluation to provide a holistic perspective on the users’ grasp of explanations (Hoffman et al., 2018; Cheng et al., 2019). In the objective evaluation, we ask users to choose, using the explanations, one neural network from three candidates that is the most accurate on unobserved data. The users are presented with several input instances, on which the three neural networks make exactly the same decisions. The only differences between the three networks lie in their explanations. Hence, to make the correct selection, the users must understand the explanations. The subjective evaluation contains 13 questions requiring users to self-report their perceptions of static explanations from three aspects: comprehension, acceptance, and trust of static explanations. The questions are adopted from existing works that have been validated in prior research.
The acceptance of explanations by end users, i.e. their intentions to use explanations, indicates the effectiveness and usability of those explanations. We measure users’ acceptance by their perceived usefulness, ease of use, and behavioral intention of explanations. These three aspects are derived from the Technology Acceptance Model (TAM) (Davis, 1989; Davis et al., 1989; Diop et al., 2019), one of the most influential and widely applied theories to explain individuals’ acceptance and use of information systems.
Trust has been recognized as an important factor in human-AI collaboration as it mediates the human’s reliance on AI models, thus directly affecting the effectiveness of the human-AI collaboration (Sebo et al., 2020; Seaborn et al., 2021; Doshi-Velez and Kim, 2017). However, the literature reports mixed effects of explanations on trust. Though several studies have shown that increased transparency in machine learning systems does not significantly impact people’s trust (Cheng et al., 2019; Schaffer et al., 2019; Kunkel et al., 2019; Poursabzi-Sangdeh et al., 2021), others find that explanations can improve users’ trust (Larasati et al., 2020; Yang et al., 2020) and can potentially lead to misplaced trust in machine learning models (Wang and Yin, 2021; Zhang et al., 2020). Given these contrasting views, we measure users’ trust in explanations to investigate the effects of conversations.
@@
From the results, we discover that free-form conversations with XAI experts in the Wizard-of-Oz setting significantly improve participants’ comprehension of static explanations and their ability to select the best AI models based on explanations. Besides, they report an increase in perceived acceptance and trust of static explanations after conversations. By analyzing the collected conversations, we observe that participants’ questions are unanticipated and diverse. This explains why static explanations do not resolve users’ confusion about neural networks and indicates the importance and necessity of conversational explanations. We also discuss the implication of our study for the future design of dialogue systems to explain static explanations. Our study underscores the effects of free-form conversations on neural network explainability in practice and provides insights into the future development of conversational explanations. To the best of our knowledge, we present the first study that investigates the effects of Wizard-of-Oz conversations on neural network explainability in practice.
@@ 2. Related Work
Explainable Artificial Intelligence (XAI) refers to those models that can explain either the learning process or the outcome of AI predictions to human users (Yang et al., 2019). In this section, we will discuss related work about static and interactive explanation methods in XAI.
@@ 2.1. Static Explanation
Static XAI is usually categorized into two groups: self-explanatory models and post-hoc methods. Post-hoc methods can be categorized into feature attribution methods and example-based methods. Self-explanatory models are inherently transparent, offering clarity in their decision-making processes and facilitating explainability (Danilevsky et al., 2020; Bodria et al., 2023). Self-explanatory models can be viewed as the opposite of black box models. Examples of such models include linear regression, logistic regression, decision trees, Naive Bayes, attention mechanism (Bahdanau et al., 2014), decision sets (Lakkaraju et al., 2016), rule-based models (Rudziński, 2016; Yang et al., 2017), among others. However, the requirements of self-explanatory models place constraints on model design, which may cause them to underperform in complex tasks. Conversely, the majority of recent XAI methods are post-hoc XAI methods, which can be used for an already developed or trained model that is usually not inherently transparent (Selvaraju et al., 2017; Ribeiro et al., 2016; Chen et al., 2021; Verma et al., 2020; Adadi and Berrada, 2018; Bodria et al., 2023). These methods often do not attempt to explain how the model works internally, but instead, employ separate techniques to extract explanatory information. Post-hoc XAI methods can be viewed as reverse engineering processes that employ other independent explanatory models or techniques to extract explanatory information without altering, elucidating, or even understanding the inner workings of the original black-box model. There are two main groups of methods to generate post-hoc XAI explanations, i.e., feature attribution methods and example-based methods.
@@ 2.1.1. Feature Attribution Methods
Example-based methods (Chen et al., 2021; Verma et al., 2020; Tran et al., 2021; Mothilal et al., 2020; Poyiadzi et al., 2020; Wijaya, 2021; Jeyakumar et al., 2020) refer to those that explain predictions of black-box models by identifying and presenting a selection of similar or representative instances. Those examples can be selected or generated from different perspectives, such as training data points that are the most influential to the parameters of a prediction model or the predictions themselves (Chen et al., 2019, 2021; Yoon et al., 2020), counterfactual examples that are similar to the input query but with different predictions (Verma et al., 2020; Wachter et al., 2017; Sharma et al., 2019; Karimi et al., 2020; Tran et al., 2021; Mothilal et al., 2020; Poyiadzi et al., 2020; Wijaya, 2021), or prototypes that contain semantically similar parts to input instances (Croce et al., 2019; Bien and Tibshirani, 2011; Jeyakumar et al., 2020; Mikolov et al., 2013; Doshi-Velez et al., 2015; Kim et al., 2016).
In this work, we mainly focus on feature attribution methods as they directly highlight the importance of input features, making the model’s decision-making process more intuitive (Kim et al., 2023). Specifically, we select Grad-CAM from gradient-based methods and LIME from surrogate methods to conduct conversational explanations with participants.
@@ 2.2. Interactive Explanation
Several studies emphasize the need for interactivity in XAI methods (Lakkaraju et al., 2022; Abdul et al., 2018; Schmid and Wrede, 2022; Rohlfing et al., 2020). For instance, Lakkaraju et al. (Lakkaraju et al., 2022) find that decision-makers strongly prefer interactive explanations. Similarly, a literature analysis by Abdul et al. (Abdul et al., 2018) suggests that interactions can help users progressively explore and gather insights from static explanations. Rohlfing et al. (Rohlfing et al., 2020) reason that explanations should be co-constructed in an interaction between the explainer and the explainee, adapting to individual differences since the human understanding process is dynamic. From an interdisciplinary perspective, Schmid and Wrede (Schmid and Wrede, 2022) underscore the necessity of user-XAI interactions to adapt to diverse information requirements.
@@
To integrate interactivity and explainability, two primary methodologies emerge. One group of methods focuses on using explanations to help users provide feedback about improving machine learning models. In these methods, the interactivity lies in the cycle of model explanation, user feedback, and model improvement. Explanations aim to help users better understand model decisions and provide valuable feedback. As a result, machine learning models can be incrementally trained with additional loss terms from explanatory feedback (Kulesza et al., 2015; Lertvittayakumjorn et al., 2020; Liang et al., 2020; Ross et al., 2017; Schramowski et al., 2020; Smith-Renner et al., 2020) or with added data points (Alkan et al., 2022; Biswas and Parikh, 2013; Teso and Kersting, 2019; Teso et al., 2021). However, these methods are aimed at machine learning practitioners who can well understand and utilize explanations. Another group focuses on enhancing user understanding of explanations by allowing them to modify the model input and observe changes in the corresponding output. Such interactivity has been shown to improve user comprehension and the perceived utility of AI models (Cheng et al., 2019; Liu et al., 2021a). For instance, Tenney et al. (Tenney et al., 2020) and Hohman et al. (Hohman et al., 2019) propose different user interfaces that allow for debugging and understanding machine learning models by examining input-output relationships. However, a rudimentary understanding of machine learning is still required for effective utilization of these interfaces, such as the generic relation between input and output, or what model properties the interpretations reveal.
@@
HCI researchers have recently proposed that XAI methods should align with the ways humans naturally explain mechanisms. Specifically, Lombrozo (Lombrozo, 2006) argues that an explanation is a byproduct of a conversational interaction process between an explainer and an explainee. Miller (Miller, 2019) argues that explanations should contain a communication process, where the explainer interactively provides the information required for the explainee to understand the causes of the event through conversations. Hilton(Hilton, 1996) suggests a two-stage process: diagnosis to identify causation and then an explanation to communicate the diagnosis. Building on this perspective of human explanations, recent works envision ”explainability as dialogue” to provide explanations suitable for a wide range of layperson users (Liao et al., 2020; Dazeley et al., 2021; Feldhus et al., 2022; Lakkaraju et al., 2022). While there is much theoretical analysis about the significance of conversations in explainability, practical investigations into their impact on users remain limited. In this study, we explore the role of conversations in enhancing users’ comprehension of static explanations, and how they affect users’ acceptance, trust, and collaboration with these explanations.

### v2

@@ Abstract.
Research in explainable AI (XAI) aims to provide insights into the decision-making process of opaque AI models. To date, most XAI methods offer one-off and static explanations, which cannot cater to the diverse backgrounds and understanding levels of users. With this paper, we investigate if free-form conversations can enhance users’ comprehension of static explanations in image classification, improve acceptance and trust in the explanation methods, and facilitate human-AI collaboration. We conduct a human-subject experiment with 120 participants. Half serve as the experimental group and engage in a conversation with a human expert regarding the static explanations, while the other half are in the control group and read the materials regarding static explanations independently. We measure the participants’ objective and self-reported comprehension, acceptance, and trust of static explanations. Results show that conversations significantly improve participants’ comprehension, acceptance (Davis, 1989), trust, and collaboration with static explanations, while reading the explanations independently does not have these effects and even decreases users’ acceptance of explanations. Our findings highlight the importance of customized model explanations in the format of free-form conversations and provide insights for the future design of conversational explanations.
@@ 1. Introduction
The rapid advancement of Artificial Intelligence (AI) is largely powered by opaque deep neural networks (DNNs), which are difficult to interpret by humans (Bodria et al., 2023). The lack of transparency prevents verification of AI decisions by human domain experts and is especially concerning in areas of high-stake decisions, such as healthcare and law enforcement, where erroneous algorithmic decisions could lead to severe consequences (Caruana et al., 2015; Cai et al., 2019; Zheng et al., 2023) and erosion of public trust (Quinn et al., 2021; Powles and Hodson, 2017). To improve the explainability of AI models, numerous eXplainable Artificial Intelligence (XAI) methods have been proposed (for detailed reviews, we refer readers to Yang et al. (2019); Danilevsky et al. (2020); Bodria et al. (2023)). It has been reported that explainability enhances user understanding (Bansal et al., 2021) and trust (González et al., 2021; Luo et al., 2022) in AI models, improves human-AI collaboration in decision-making (Nguyen et al., 2022; Lai and Tan, 2019), and helps AI developers identify and rectify model errors (Adebayo et al., 2020; Idahl et al., 2021). Despite these successes, a number of recent studies find that the explanations often do not resolve user confusion regarding the neural networks they are purported to explain (Poursabzi-Sangdeh et al., 2021; Zhang et al., 2020; Bansal et al., 2021; Wang and Yin, 2021; Lakkaraju et al., 2022; Liao et al., 2020; Slack et al., 2023; Shen et al., 2023). These seemingly conflicting findings warrant further investigation.
@@
We postulate that two major factors contribute to the ineffectiveness of AI explanations. First, the explanations do not properly account for average users’ knowledge of machine learning, which may be insufficient to establish causal relations between the explanations and the model behaviors (Poursabzi-Sangdeh et al., 2021; Ma et al., 2023; Springer and Whittaker, 2019; He et al., 2023). Communication theory posits that effective communication requires the senders and receivers to establish common ground (Clark and Marshall, 1981; Clark and Brennan, 1991). However, experts usually find it hard to accurately estimate what laypeople know (Wittwer et al., 2008; Wilkesmann and Wilkesmann, 2011; Miller, 2019). To make matters worse, underestimating and overestimating the receivers’ knowledge level are equally detrimental to communication (Wittwer et al., 2008; Lakkaraju et al., 2022). As a result, the explanations designed by experts are almost always at a mismatch with the laypersons’ actual knowledge level.
@@
Second, users of XAI have diverse intentions and information needs (Wang and Yin, 2021; Ehsan et al., 2024; Liao et al., 2020; He et al., 2023). For example, Liao and Varshney (2021) identifies five different objectives of users of explanations, including model debugging, assessing the capabilities of AI systems, making informed decisions, seeking recourse or contesting the AI, and auditing for legal or ethical compliance. One static explanation usually cannot satisfy all objectives and purposes. Therefore, researchers have suggested injecting interactivity to model explanations in order to establish common ground, address knowledge gaps, and create customized explanations that adapt to the users (Lakkaraju et al., 2022; Abdul et al., 2018; Schmid and Wrede, 2022; Rohlfing et al., 2020; Cheng et al., 2019; Guesmi et al., 2023).
@@
Existing work on interactive explanations can be broadly categorized into two types. The first type, interactive machine learning (Amershi et al., 2014; Fails and Olsen Jr, 2003), allows users to provide feedback and suggestions to the machine learning model using model explanations. Their primary goal is to improve machine learning performances, rather than explaining model behaviors to layperson users. In this setting, explanations have been shown to improve user satisfaction (Smith-Renner et al., 2020) and feedback quality (Liang et al., 2020; Kulesza et al., 2015). The second type aims to elucidate model behaviors by allowing users to freely modify input features and observe how outputs change while showing feature attribution explanations (Cheng et al., 2019; Liu et al., 2021a; Tenney et al., 2020; Hohman et al., 2019). This type of interactivity has been shown to improve user understanding (Cheng et al., 2019) and perceived usefulness (Liu et al., 2021a) of AI models. However, the effective use of these interactive approaches still requires a rudimentary understanding of machine learning, such as the generic relation between input and output, or what model properties the interpretations reveal. These interactive explanations cannot answer most types of follow-up questions laypeople may have.
@@
Free-form conversations that accompany static explanations are arguably the most versatile mode of interaction as they allow users to ask arbitrary follow-up questions and receive explanations tailored to their backgrounds and needs (Liao et al., 2020; Feldhus et al., 2022; Lakkaraju et al., 2022). Through interviews with decision-makers, Lakkaraju et al. (2022) discover that they have a strong preference for explanations in natural language dialogue. They argue that conversational explanations satisfy five requirements of interactive explanations and are ideal for users with limited machine learning knowledge. With the progress in conversational characters (Zhang et al., 2022; Ni et al., 2023; Shuster et al., 2022), especially knowledge-based question answering (Lan et al., 2021; Luo et al., 2023; Zhang et al., 2023) powered by large language models (Touvron et al., 2023; Ouyang et al., 2022; Zhao et al., 2023), AI systems that can answer questions about their own decisions appear to be within our reach in the near future. However, before investing effort to develop such a chatbot, it would be beneficial to empirically quantify the effects of conversational explanations.
@@
In the current study, we conduct Wizard-of-Oz experiments to investigate how conversations assist users in understanding static explanations of image classification models, improving acceptance and trust in XAI methods, and selecting the best AI models based on explanations. Specifically, a total of 120 participants join our experiments. We first present them with static explanations for an image classification task and measure their objective understanding and subjective perceptions of static explanations. After that, half of the participants, who are assigned to the experimental group, seek to clarify any doubts with an online textual conversation with an AI system, played by human XAI experts. The other half of the participants, assigned to the control group, read materials about the static explanations independently. After the conversation or reading session, participants complete the same pre-session measurements. From the results, we estimate the effects of conversational explanations.
@@
The experimental measurements include both an objective component and a subject component of the users’ understanding and perception. In the objective evaluation, from three candidate neural networks, the users need to choose one network that would be the most accurate on test data so far unobserved, using information from the static explanations. This task, known as model selection, is one of the most fundamental tasks for machine learning practitioners (Anderson and Burnham, 2004). By design, the three candidate networks make exactly the same predictions on the same inputs but have different rationales for the predictions, as revealed by the static explanations. Hence, the only way for the users to make the right choice is to correctly understand the explanations. The subjective evaluation contains 13 questions requiring users to self-report three aspects of their perceptions of the static explanations: comprehension, acceptance, and trust.
@@
Results show that free-form conversations with XAI experts in the Wizard-of-Oz setting significantly improve comprehension, acceptance, trust, and collaboration with static explanations. Our study underscores the effects of free-form conversations on neural network explainability in practice and provides insights into the future development of conversational explanations. To the best of our knowledge, this is the first study of how free-form conversations may facilitate neural network explainability in practice.
@@ 2. Related Work
In this section, we review three bodies of research that motivate our study. First, we explore the existing work of static Explainable Artificial Intelligence (XAI). Second, we discuss interactive explanations, especially the limitations of existing methods and the need for conversations to enhance explainability. Lastly, we examine different types of human-AI collaboration and the design of the subjective evaluation during collaboration.
@@ 2.1. Static Explanation
Explainable Artificial Intelligence (XAI) refers to those models that can explain either the learning process or the outcome of AI predictions to human users (Yang et al., 2019). Static XAI involves models that provide a fixed, one-time explanation, without the capability for further user interaction or exploration. They are usually categorized into two groups: self-explanatory models and post-hoc methods. Post-hoc methods can be categorized into feature attribution methods and example-based methods. Self-explanatory models are inherently transparent, offering clarity in their decision-making processes and facilitating explainability (Danilevsky et al., 2020; Bodria et al., 2023). Examples of such models include linear regression, logistic regression, decision trees, Naive Bayes, attention mechanism (Bahdanau et al., 2015), decision sets (Lakkaraju et al., 2016), rule-based models (Rudziński, 2016; Yang et al., 2017a), among others. However, the requirements of self-explanatory models place constraints on model design, which may cause them to underperform in complex tasks. Conversely, the majority of recent XAI methods are post-hoc XAI methods, which can be used for an already developed model that is usually not inherently transparent (Selvaraju et al., 2017; Ribeiro et al., 2016; Chen et al., 2021; Verma et al., 2020; Adadi and Berrada, 2018; Bodria et al., 2023). These methods often do not attempt to explain how the model works internally, but instead, employ separate techniques to extract explanatory information. Post-hoc XAI methods can be viewed as reverse engineering processes that employ other independent explanatory models or techniques to extract explanatory information without altering, elucidating, or even understanding the inner workings of the original black-box model. There are two main groups of methods to generate post-hoc XAI explanations, i.e., feature attribution methods and example-based methods.
@@ 2.1.1. Feature Attribution Methods
Example-based methods (Chen et al., 2021; Verma et al., 2020; Tran et al., 2021; Mothilal et al., 2020; Poyiadzi et al., 2020; Jeyakumar et al., 2020) refer to those that explain predictions of black-box models by identifying and presenting a selection of similar or representative instances. Those examples can be selected or generated from different perspectives, such as training data points that are the most influential to the parameters of a prediction model or the predictions themselves (Chen et al., 2019, 2021; Yoon et al., 2020), counterfactual examples that are similar to the input query but with different predictions (Verma et al., 2020; Wachter et al., 2017; Sharma et al., 2020; Karimi et al., 2020; Tran et al., 2021; Mothilal et al., 2020; Poyiadzi et al., 2020), or prototypes that contain semantically similar parts to input instances (Croce et al., 2019; Bien and Tibshirani, 2011; Jeyakumar et al., 2020; Mikolov et al., 2013; Doshi-Velez et al., 2015; Kim et al., 2016).
In this work, we mainly focus on feature attribution methods as they directly highlight the importance of input features, making the decision-making process of models more intuitive (Kim et al., 2023) than example-based methods for laypeople. Specifically, we select Grad-CAM from gradient-based methods and LIME from surrogate methods to conduct conversational explanations with participants.
@@ 2.2. Interactive Explanation
Several studies emphasize the need for interactivity in XAI methods (Lakkaraju et al., 2022; Abdul et al., 2018; Schmid and Wrede, 2022; Rohlfing et al., 2020). For instance, Lakkaraju et al. (2022) find that decision-makers strongly prefer interactive explanations. Similarly, a literature analysis by Abdul et al. (2018) suggests that interactions can help users progressively explore and gather insights from static explanations. Rohlfing et al. (2020) reason that explanations should be co-constructed in an interaction between the explainer and the explainee, adapting to individual differences since the human understanding process is dynamic. From an interdisciplinary perspective, Schmid and Wrede (2022) underscore the necessity of user-XAI interactions to adapt to diverse information requirements.
@@
To integrate interactivity and explainability, two primary methodologies emerge. One group of methods focuses on using explanations to help users provide feedback about improving machine learning models. In these methods, the interactivity lies in the cycle of model explanation, user feedback, and model improvement. Explanations aim to help users better understand model decisions and provide valuable feedback. As a result, machine learning models can be incrementally trained with additional loss terms from explanatory feedback (Kulesza et al., 2015; Lertvittayakumjorn et al., 2020; Liang et al., 2020; Ross et al., 2017; Schramowski et al., 2020; Smith-Renner et al., 2020) or with added data points (Alkan et al., 2022; Biswas and Parikh, 2013; Teso and Kersting, 2019; Teso et al., 2021). However, these methods are aimed at machine learning practitioners who can well understand and utilize explanations. Another group focuses on enhancing user understanding of explanations by allowing them to modify the model input and observe changes in the corresponding output. Such interactivity has been shown to improve user comprehension and the perceived utility of AI models (Cheng et al., 2019; Liu et al., 2021a). For instance, Tenney et al. (2020) and Hohman et al. (2019) propose different user interfaces that allow for debugging and understanding machine learning models by examining input-output relationships. However, a rudimentary understanding of machine learning is still required for effective utilization of these interfaces, such as the generic relation between input and output, or what model properties the interpretations reveal.
@@
HCI researchers have recently proposed that XAI methods should align with the ways humans naturally explain mechanisms. Specifically, Lombrozo (2006) argues that an explanation is a byproduct of a conversational interaction process between an explainer and an explainee. Miller (2019) argues that explanations should contain a communication process, where the explainer interactively provides the information required for the explainee to understand the causes of the event through conversations. Building on this perspective of human explanations, recent works envision ”explainability as dialogue” to provide explanations suitable for a wide range of layperson users (Liao et al., 2020; Feldhus et al., 2022; Lakkaraju et al., 2022). While there is much theoretical analysis about the significance of conversations in explainability, practical investigations into their impact on users remain limited. In this context, two previous works have investigated the practical effect of conversations for explainability (Slack et al., 2023; Shen et al., 2023). Shen et al. (2023) apply conversational explanations to scientific writing tasks, observing improvements in productivity and sentence quality. Slack et al. (2023) design dialogue systems to help users better understand machine learning models on diabetes prediction, rearrest prediction, and loan default prediction tasks. Despite these advances, the conversations in these studies are generated based on templates and cope with limited predefined user intentions. In this study, we explore the role of free-form conversations in enhancing users’ comprehension of static explanations, and how they affect users’ acceptance, trust, and collaboration with these explanations.